Shift report 20010813 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Mon Aug 13 2001 - 04:12:17 EDT

  • Next message: Apache: "Shift report 20010813 08:00-16:00"

    ------------------------------------------------------------
    Date: Mon Aug 13 04:12:16 2001, Shift: 00:00-08:00
    Supervisor: YK Lee, with: Michael Murray
    Log book#7 pages 70 - 70
    
    Shift Summary:
    ------------------------------------------------------------
    No beam, problems with DX vacuum and quenching
    No achivements.
    D2 oscilated +-30A about its set-point of 922A. The operator
    at MCR did not know who to call. Since beam is not expected
    before 8am we decided to turn off this magnet and go home.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Mon Aug 13 2001 - 04:12:47 EDT